OpenStudy (anonymous):

what is the period and amplitude of y=-1/2cos(x/3)

OpenStudy (anonymous):

already answered???!!!

OpenStudy (anonymous):

was it really?

OpenStudy (anonymous):

it was not

OpenStudy (lalaly):

amplitude = 1/2 period = 2pi/(1/3) = 6pi